A study found the following data for the number of intersections in a certain neighborhood:

1 8 6 1 1 5 3 0 0 4

0 0 1 2 1 4 0 4 0 3

1 1 0 1 3 2 4 6 6 0

1 8 3 3 5 0 5 2 3 1

construct a histogram for the data. what proportion of these data points had at least 4 intersections? how many had at most 2 intersections?
Answer #1

Ans:

Frequency Table 20 19 Class 0-1 2-3 4-5 6-7 19 15 2 10 Your Histogram Mean 2.475 2.29813 0.80767 0.18299 Standard Deviation (

Proportion of data had at least 4 intersections=(7+3)/40=10/40=0.25

Proportion of data had at most 2 intersections=19/40=0.475